# Light-quark mass dependence of the $\Lambda(1405)$ resonance

### v1

@@ Abstract
We present the light-quark mass dependence of the $\Lambda(1405)$ resonance at leading order in a renormalizable framework of covariant chiral effective field theory. The meson-baryon scattering amplitudes, which are obtained by solving the scattering equation within time-ordered perturbation theory, follow the quark mass trajectory of the Coordinated Lattice Simulations consortium. At $M_{\pi}\approx 200$ MeV and $M_{K}\approx 487$ MeV, our parameter-free prediction of $\Lambda(1405)$ poles is consistent with the recent lattice results of BaSc Collaboration [Phys. Rev. Lett. 132, 051901 (2024)]. Varying the pion mass from $135$ MeV to $400$ MeV, we present the evolution of double-pole positions of $\Lambda(1405)$: the higher pole remains a resonance around the $\bar{K}N$ threshold; whereas the lower pole undergoes a transition from resonance to a virtual state, and ultimately to a bound state of the $\pi\Sigma$ system, which could serve as a prediction of the forthcoming lattice QCD simulations.
@@ 1 Introduction
That is one purpose of the current work to examine our renormalizable approach for the meson-baryon scattering, which has been proposed in Ref. [32]. Starting with the covariant chiral effective Lagrangians and employing the time-ordered perturbation theory (TOPT), we define an effective potential as the sum of two-particle irreducible contributions of time-ordered diagrams. The corresponding (coupled-channel) scattering equation $T=V+VGT$ can be derived self-consistently in TOPT. At leading order (LO), we take into account the full off-shell dependence in solving the scattering equation. The renormalized scattering amplitude is obtained by utilizing the subtractive renormalization scheme [33]. Note that the subtracted terms are fixed by the chiral symmetry, which is different from the often-used strategies in CUA, where the finite-cutoff or the subtraction constant varies to match the data. In this sense, our framework at LO provides a model-independent prediction for the meson-baron scattering.
As shown in Ref. [32], our approach has been successfully applied to the pion-nucleon system. Then, we extend this framework to the strangeness $S=-1$ sector with isospin $I=0$ and study the $\Lambda(1405)$ resonance at the physical point [34]. We solve the scattering equation with the four coupled channels $\pi\Sigma$, $\bar{K}N$, $\eta\Lambda$, and $K\Xi$, and obtain the renormalized $T$-matrix in the isospin limit. Two $\Lambda(1405)$ poles are found on the second Riemann sheet: the lower pole ($79.11337.7-i$ 79.1 MeV) is close to the $\pi\Sigma$ threshold, the higher pole ($8.01430.9-i$ 8.0 MeV) just below $\bar{K}N$ threshold. Our LO results are consistent with the ones of the next-to-leading order study in CUA [35].
Coming to the unphysical quark mass region, it is interesting to extrapolate our parameter-free prediction of $\Lambda(1405)$ poles and compare it with the BaSc lattice data. If such an agreement is reached, it indicates that our approach exhibits the predictive capability to some extent. Therefore, it enables us to investigate the quark mass dependence of the $\Lambda(1405)$ state. Similar studies along this line can be found in Refs. [36, 37, 38, 39]. Here we focus on the results along with the quark-mass trajectory of the CLS ensembles. Our model variables, the pseudoscalar meson masses, octet baryon masses, vector meson masses, and meson decay constants, have been calculated in Refs. [40, 41, 42] based on the CLS configuration. It provides a good opportunity to investigate the light-quark mass dependence of $\Lambda(1405)$, which can be directly verified by the forthcoming results of the BaSc Collaboration.
@@ 2.1 Meson-baryon scattering in TOPT
For self-completeness, here we briefly lay out the formalism to study the $S=-1$ meson-baryon scattering in the SU(3) unitarized chiral EFT based on a renormalizable framework. The details can be found in Ref. [34].
At LO, the meson-baryon scattering potential contains the vector-meson exchange contribution, the Born and crossed-Born terms. Note that we employ the vector-meson exchange to saturate the often-used Weinberg-Tomozawa term since the former has a better ultraviolet behavior without changing the low-energy physics. For the strangeness $S=-1$ sector with isospin $I=0$, the LO interaction for the process $M_{i}(q_{1})+B_{i}(q_{2})\to M_{j}(q_{2})+B_{j}(p_{2})$ is given as in TOPT {strip}
$V_{\mathrm{LO}}&=-\frac{1}{32F_{0}^{2}}\sum\limits_{V}C_{M_{j}B_{j},M_{i}B_{i}}^{V}\frac{M_{V}^{2}}{\omega_{V}(\bm{q}_{1}-\bm{q}_{2})}\left(\omega_{M_{i}}(\bm{q}_{1})+\omega_{M_{j}}(\bm{q}_{2})\right)\\ &\quad\times\Biggl{[}\frac{1}{E-\omega_{B_{i}}(\bm{p}_{1})-\omega_{V}(\bm{q}_{1}-\bm{q}_{2})-\omega_{M_{j}}(\bm{q}_{2})}+\frac{1}{E-\omega_{B_{j}}(\bm{p}_{2})-\omega_{V}(\bm{q}_{1}-\bm{q}_{2})-\omega_{M_{i}}(\bm{q}_{1})}\Biggr{]},\\ &+\frac{1}{4F_{0}^{2}}\sum\limits_{B}C_{M_{j}B_{j},M_{i}B_{i}}^{B}\frac{m_{B}}{\omega_{B}(\bm{P})}\frac{(\bm{\sigma}\cdot\bm{q}_{2})(\bm{\sigma}\cdot\bm{q}_{1})}{E-\omega_{B}(\bm{P})}\\ &+\frac{1}{4F_{0}^{2}}\sum\limits_{B}\tilde{C}_{M_{j}B_{j},M_{i}B_{i}}^{B}\frac{m_{B}}{\omega_{B}(\bm{K})}\frac{(\bm{\sigma}\cdot\bm{q}_{1})(\bm{\sigma}\cdot\bm{q}_{2})}{E-\omega_{M_{i}}(\bm{q}_{1})-\omega_{M_{j}}(\bm{q}_{2})-\omega_{B}(\bm{K})},\end{aligned}$ (1)
where $F_{0}$ and $M_{V}$ are the pseudoscalar decay constant and the vector meson mass, respectively, in the chiral limit. Both are related via the Kawarabayashi-Suzuki-Riazuddin-Fayyazuddin relation [43, 44, 45]: $M_{V}^{2}=2g^{2}F_{0}^{2}$ with the coupling $g$ of the vector-field self-interaction. The constant coefficients $C^{V}$ of the vector-meson exchange contribution are given in Table 1 of Ref. [34], and the coefficients of the Born ($C^{B}$) and crossed-Born ($\tilde{C}^{B}$) terms, as the functions of the axial vector couplings $D$ and $F$, are given in Tables 2 and 3 of Ref. [34]. In the center-of-mass (c.m.) frame, the four momenta of initial and final states are
@@
with the relative momenta $\bm{p}$, $\bm{p}^{\prime}$, and the on-shell energy $\omega_{X}(\bm{p})\equiv\sqrt{m_{X}^{2}+\bm{p}^{2}}$. Furthermore, $E$ is the total energy of the meson-baryon system, and $P=q_{1}+p_{1}=q_{2}+p_{2}$, $K=p_{1}-q_{2}=p_{2}-q_{1}$.
@@
where $i,\leavevmode\nobreak\ j,\leavevmode\nobreak\ n$ denote initial, final and intermediate particle channels of meson and baryon states. The two-body Green function, which is obtained according to the diagrammatic rules of TOPT [46], has the following form
@@
where $V_{R}$ is just the Born term and $V_{I}$ includes the remaining terms. Then, one can rewrite the scattering equation (Eq. (3)), $T=V+VGT$, as the three coupled equations:
@@
It is found that the irreducible part $T_{I}$ is finite in the removed regulator limit. The only divergent term in the $T$-matrix originates from the reducible part, $T_{R}$. One can apply the subtractive renormalization [33] to systematically remove the divergences, by replacing the meson-baryon Green function $G_{n}(E)$ with the subtracted one
$G_{n}^{S}(E)=G_{n}(E)-G_{n}(m_{B_{n}}).$ (7)
Note that such subtraction, which is an analogy to the extended-on-mass-shell (EOMS) scheme [47], corresponds to the expansion around the threshold and can systematically remove the chiral power-counting breaking terms in the iterated amplitudes. Finally, we obtain a renormalized finite $T$-matrix after solving the three coupled equations in Eq. (6).
@@ 2.2 Quark mass dependence of our model variables
For the meson-baryon scattering at LO, the most relevant quark mass dependence comes from the pseudoscalar meson masses $M_{P}$ and the octet baryon masses $m_{B}$, which appear in the chiral potential and the two-body Green function. Confronting the recent lattice results of $\Lambda(1405)$ [30, 31], we conduct the chiral analysis of the meson and baryon masses from the RQCD Collaboration [40] based on the CLS configuration, which is also utilized by the BaSc Collaboration.
@@
representing the electrically neutral isospin-averaged results [49]. The best fitting result is shown in the left panel of Fig. 1 with $a=0.25189$ GeV2 and $b=-0.42143$. Then, the quark mass dependence of $M_{\eta}$ is determined via the Gell-Mann-Okubo relation $M_{\eta}^{2}=(4M_{K}^{2}-M_{\pi})/3$.
Refer to captionRefer to caption
@@
Besides, we also require the vector meson mass $M_{V}$ and the decay constant $F_{0}$ in the chiral limit to obtain the LO potential (Eq. (1)). In practice, we use the experimental masses of vector mesons $M_{\rho,\leavevmode\nobreak\ \omega,\leavevmode\nobreak\ K^{*},\leavevmode\nobreak\ \phi}$ and the pion decay constant $F_{\pi}$ or the SU(3) averaged decay constant in Ref. [34]. The introduced difference is of higher order. As to the light-quark mass dependence of $M_{V}$, only rho meson mass $M_{\rho}$ is simulated with the quark mass trajectory $\overline{m}=m_{\mathrm{symm}}$ of the CLS configuration by Mainz lattice group [41]. To determine the quark mass dependence of the vector meson masses, we use the results of LO ChPT in the SU(3) sector [50],
@@ 3 Results and discussion
Finally, we would like to remark that the qualitative analysis of the evolution of the $\Lambda(1405)$ pole positions at different quark masses has been reported in the $\pi\Sigma$-$\bar{K}N$ coupled-channel study [38] by using the Weinberg-Tomozawa term within the traditional CUA, where the light-quark mass dependence of the decay constant and the adjustable subtraction constants are neglected. In contrast, our results of the quark mass dependence of $\Lambda(1405)$, shown in Fig. 4, could serve as a parameter-free prediction for the forthcoming lattice calculation of BaSc Collaboration. This is because we include the pion mass dependence of all the model variables, which leads to the renormalized scattering amplitudes following the quark mass trajectory of the CLS configuration.

### v2

@@ Abstract
We present the light-quark mass dependence of the $\Lambda(1405)$ resonance at leading order in a renormalizable framework of covariant chiral effective field theory. The meson-baryon scattering amplitudes, which are obtained by solving the scattering equation within time-ordered perturbation theory, follow the quark mass trajectory of the Coordinated Lattice Simulations consortium. At $M_{\pi}\approx 200$ MeV and $M_{K}\approx 487$ MeV, our parameter-free prediction of $\Lambda(1405)$ poles is consistent with the recent lattice results of BaSc Collaboration [Phys. Rev. Lett. 132, 051901 (2024)]. Varying the pion mass from $135$ MeV to $400$ MeV, we present the evolution of double-pole positions of $\Lambda(1405)$: the higher pole remains a resonance around the $\bar{K}N$ threshold; whereas the lower pole undergoes a transition from resonance to a virtual state, and ultimately to a bound state of the $\pi\Sigma$ system, which could be verified by the forthcoming lattice QCD simulations.
@@ 1 Introduction
That is one purpose of the current work: to examine our renormalizable approach for the meson-baryon scattering, which was proposed in Ref. [32]. Starting with the covariant chiral effective Lagrangians and employing the time-ordered perturbation theory (TOPT), we define an effective potential as the sum of two-particle irreducible contributions of time-ordered diagrams. The corresponding (coupled-channel) scattering equation $T=V+VGT$ can be derived self-consistently in TOPT. At leading order (LO), we take into account the full off-shell dependence in solving the scattering equation. The renormalized scattering amplitude is obtained by utilizing the subtractive renormalization scheme [33]. Note that the subtracted terms are fixed by the chiral symmetry, which is different from the often-used strategies in CUA, where the finite-cutoff or the subtraction constant varies to match the data. In this sense, our framework at LO provides a model-independent prediction for the meson-baryon scattering.
As shown in Ref. [32], our approach has been successfully applied to the pion-nucleon system. Then, we extend this framework to the strangeness $S=-1$ sector with isospin $I=0$ and study the $\Lambda(1405)$ resonance at the physical point [34]. We solve the scattering equation with the four coupled channels $\pi\Sigma$, $\bar{K}N$, $\eta\Lambda$, and $K\Xi$, and obtain the renormalized $T$-matrix in the isospin limit. Two $\Lambda(1405)$ poles are found on the second Riemann sheet: the lower pole ($79.11337.7-i$ 79.1 MeV) is close to the $\pi\Sigma$ threshold, the higher pole ($8.01430.9-i$ 8.0 MeV) just below $\bar{K}N$ threshold. Our LO results are consistent with the ones of the next-to-leading order (NLO) study in CUA [35].
Coming to the unphysical quark mass region, it is interesting to extrapolate our parameter-free prediction of $\Lambda(1405)$ poles and compare it with the BaSc lattice data. If such an agreement is reached, it indicates that our approach exhibits the predictive capability to some extent. Therefore, it enables us to investigate the quark mass dependence of the $\Lambda(1405)$ state. Similar studies along this line can be found in Refs. [36, 37, 38, 39, 40, 41]. Here we focus on the results along with the quark-mass trajectory of the CLS ensembles. Our model variables, the pseudoscalar meson masses, octet baryon masses, vector meson masses, and meson decay constants, have been calculated in Refs. [42, 43, 44] based on the CLS configuration. It provides a good opportunity to investigate the light-quark mass dependence of $\Lambda(1405)$, which could be verified by the forthcoming results of the BaSc Collaboration. Note that this LO study can be thought as a first step and the higher order corrections are planned in the future to refine our conclusion.
@@ 2 Formalism
In this section, we first lay out the formalism to study the $S=-1$ meson-baryon scattering in the SU(3) unitarized chiral EFT based on a renormalizable framework. The details can be found in Ref. [34]. After that, we investigate the light-quark mass dependence of our model variables by the combination of chiral formulae and lattice QCD data based on the CLS configurations in order to study the evolution of $\Lambda(1405)$ poles. It is worth noting that the three-body effects from $\pi\pi\Lambda$ states could be involved in the $\Lambda(1405)$ region, particularly as we move closer to the physical point. This complicated estimation is beyond the scope of this Letter, and thus we do not include it, as is usually done in the literature. Interested readers can refer to e.g. Refs. [45, 46] for more details about the three-body scatterings.
@@ 2.1 Meson-baryon scattering in TOPT
At LO, the meson-baryon scattering potential contains the vector-meson exchange contribution, the Born and crossed-Born terms. Note that we employ the vector-meson exchange (VME) to saturate the often-used Weinberg-Tomozawa (WT) term since the former has a better ultraviolet behavior without changing the low-energy physics. For the strangeness $S=-1$ sector with isospin $I=0$, the LO interaction for the process $M_{i}(q_{1})+B_{i}(q_{2})\to M_{j}(q_{2})+B_{j}(p_{2})$ is given as in TOPT {strip}
$V_{\mathrm{LO}}&=-\frac{1}{32F_{0}^{2}}\sum\limits_{V}C_{M_{j}B_{j},M_{i}B_{i}}^{V}\frac{M_{V}^{2}}{\omega_{V}(\bm{q}_{1}-\bm{q}_{2})}\left(\omega_{M_{i}}(\bm{q}_{1})+\omega_{M_{j}}(\bm{q}_{2})\right)\\ &\quad\times\Biggl{[}\frac{1}{E-\omega_{B_{i}}(\bm{p}_{1})-\omega_{V}(\bm{q}_{1}-\bm{q}_{2})-\omega_{M_{j}}(\bm{q}_{2})}+\frac{1}{E-\omega_{B_{j}}(\bm{p}_{2})-\omega_{V}(\bm{q}_{1}-\bm{q}_{2})-\omega_{M_{i}}(\bm{q}_{1})}\Biggr{]}\\ &+\frac{1}{4F_{0}^{2}}\sum\limits_{B}C_{M_{j}B_{j},M_{i}B_{i}}^{B}\frac{m_{B}}{\omega_{B}(\bm{P})}\frac{(\bm{\sigma}\cdot\bm{q}_{2})(\bm{\sigma}\cdot\bm{q}_{1})}{E-\omega_{B}(\bm{P})}\\ &+\frac{1}{4F_{0}^{2}}\sum\limits_{B}\tilde{C}_{M_{j}B_{j},M_{i}B_{i}}^{B}\frac{m_{B}}{\omega_{B}(\bm{K})}\frac{(\bm{\sigma}\cdot\bm{q}_{1})(\bm{\sigma}\cdot\bm{q}_{2})}{E-\omega_{M_{i}}(\bm{q}_{1})-\omega_{M_{j}}(\bm{q}_{2})-\omega_{B}(\bm{K})},\end{aligned}$ (1)
where $F_{0}$ and $M_{V}$ are the pseudoscalar decay constant and the vector meson mass, respectively, in the chiral limit. Both are related via the Kawarabayashi-Suzuki-Riazuddin-Fayyazuddin relation [47, 48, 49]: $M_{V}^{2}=2g^{2}F_{0}^{2}$ with the coupling $g$ of the vector-field self-interaction. The constant coefficients $C^{V}$ of the VME contribution are given in Table 1 of Ref. [34], and the coefficients of the Born ($C^{B}$) and crossed-Born ($\tilde{C}^{B}$) terms, as the functions of the axial vector couplings $D$ and $F$, are given in Tables 2 and 3 of Ref. [34]. In the center-of-mass (c.m.) frame, the four momenta of initial and final states are
@@
with the relative momenta $\bm{p}$, $\bm{p}^{\prime}$, and the on-shell energy $\omega_{X}(\bm{p})\equiv\sqrt{m_{X}^{2}+\bm{p}^{2}}$. Furthermore, $E$ is the total energy of the meson-baryon system, and $P=q_{1}+p_{1}=q_{2}+p_{2}$, $K=p_{1}-q_{2}=p_{2}-q_{1}$. Note that the crossed Born term introduces poles when iterated in the scattering equation. To avoid this technical complication, as done in Refs. [50, 51], we take the total energy $E$ in the denominator as the lowest threshold of the relevant channels. Such approximation does not affect the final results because the contribution of the crossed Born term is rather small to the $S$-wave meson-baryon scattering.
@@
where $i,\leavevmode\nobreak\ j,\leavevmode\nobreak\ n$ denote the initial, final and intermediate particle channels of meson and baryon states. The two-body Green function, which is obtained according to the diagrammatic rules of TOPT [50], has the following form
@@
where $V_{R}$ is the Born term, which can separate into two diagrams by cutting only the single-baryon line, and $V_{I}$ includes the remaining terms, which cannot do such separation. Then, one can rewrite the scattering equation (Eq. (3)), $T=V+VGT$, as the three coupled equations:
@@
It is found that the irreducible part $T_{I}$ is finite in the removed regulator limit. To demonstrate this fact, we take the vector-meson-exchange potential as an example. First, we investigate the UV behavior of the once-iterated contribution $VGV$ at one-loop level:
$I_{VGV}=\int\frac{d^{3}k}{(2\pi)^{3}}V_{\mathrm{VME}}(p^{\prime},k)\,G(k)\,V_{\mathrm{VME}}(k,p),$ (7)
When the integral momentum $k$ goes to infinity, the one-loop integral becomes
$I_{VGV}\to\int\frac{d^{3}k}{(2\pi)^{3}}\frac{1}{k}\,\frac{1}{k^{3}}\,\frac{1}{k},$ (8)
which converges. This benefit arises from the VME potential, which has a milder UV behavior compared to the often-used WT term. One can easily verify that the infinite iteration of $V_{\mathrm{VME}}$ in the ladder diagram also converges. Besides, the crossed-Born term presents a similar UV behavior to the VME counterpart. Because the half-off-shell potential $V_{\mathrm{CB}}(k,p)$ approaches
$V_{\mathrm{CB}}(k,p)\to\frac{\bm{\sigma\cdot\bm{p}}\,\bm{\sigma}\cdot{\bm{k}}}{k^{2}},$ (9)
when the integral momentum $k$ goes to infinity.
Regarding the reducible part, because the iteration of the Born term is quadratically divergent, thus the only divergent term in the total $T$-matrix originates from the reducible part $T_{R}$. To achieve renormalization, we first rewrite the Born potential in a separable form,
$V_{R}(p^{\prime},p;E)=\xi^{T}(p^{\prime})\,C(E)\,\xi(p),$ (10)
where $\xi^{T}(p)$ is defined as $\xi^{T}(p)=(1,\,p)$, and $C(E)$ denotes a $2\times 2$ matrix. Then, the reducible $T_{R}$ matrix can be expressed in a separable form
$T_{R}(p^{\prime},p;E)=\xi^{T}(p^{\prime})\,\chi(E)\,\xi(p),$ (11)
where $\chi(E)$, as given below, is divergent.
$\chi(E)=\left[C(E)^{-1}-\xi\,G(E)\,\xi^{T}-\xi\,G(E)\,T_{I}^{S}\,G(E)\,\xi^{T}\right]^{-1}.$ (12)
Then, one can apply the subtractive renormalization [33] by replacing the meson-baryon Green function $G(E)$ with the subtracted one
$G^{S}(E)=G(E)-G(m_{B}),$ (13)
to systematically remove the divergences. This corresponds to taking into account the counter-terms generated by the renormalization of the baryon masses and the meson-baryon coupling constants. Note that such subtraction, which is an analogy to the extended-on-mass-shell (EOMS) scheme [52], corresponds to the expansion around the threshold and can systematically remove the chiral power-counting breaking terms in the iterated amplitudes. Finally, we obtain a renormalized $T$-matrix
$T=T_{I}^{S}+\left(\xi^{T}+T_{I}^{S}\,G^{S}\,\xi^{T}\right)\,\chi^{S}(E)\,\left(\xi+\xi\,G^{S}\,T_{I}^{S}\right),$ (14)
where the superscript $S$ demonstrates the replacement of $G$ as $G^{S}$ in the expressions of $T_{I}$ and $\chi(E)$.
@@ 2.2 Quark mass dependence of our model variables
First, we investigate the quark mass dependence of the pseudoscalar meson masses $M_{P}$ and the octet baryon masses $m_{B}$, which appear in both of the LO chiral potential and the two-body Green function. Confronting the recent lattice results of $\Lambda(1405)$ [30, 31], we conduct the chiral analysis of the meson and baryon masses from the RQCD Collaboration [42] based on the CLS configuration, which is also utilized by the BaSc Collaboration.
@@
representing the electrically neutral isospin-averaged results [54]. The best fitting result is shown in the left panel of Fig. 1 with $a=0.25189$ GeV2 and $b=-0.42143$. Then, the quark mass dependence of $M_{\eta}$ is determined via the Gell-Mann-Okubo relation $M_{\eta}^{2}=(4M_{K}^{2}-M_{\pi}^{2})/3$.
@@
Besides, we also require the vector meson mass $M_{V}$ and the decay constant $F_{0}$ in the chiral limit to obtain the LO potential (Eq. (1)). In practice, we use the experimental masses of vector mesons $M_{\rho,\leavevmode\nobreak\ \omega,\leavevmode\nobreak\ K^{*},\leavevmode\nobreak\ \phi}$ and the pion decay constant $F_{\pi}$ or the SU(3) averaged decay constant in Ref. [34]. The introduced difference between the chiral limit values and the physical values of the the vector meson masses and the meson decay constant is of higher order. As to the light-quark mass dependence of $M_{V}$, only rho meson mass $M_{\rho}$ is simulated with the quark mass trajectory $\overline{m}=m_{\mathrm{symm}}$ of the CLS configuration by Mainz lattice group [43]. To determine the quark mass dependence of the vector meson masses, we use the results of LO ChPT in the SU(3) sector [55],
@@ 3 Results and discussion
We would like to remark that the similar analysis of the evolution of the $\Lambda(1405)$ pole positions at different quark masses has been reported in the $\pi\Sigma$-$\bar{K}N$ coupled-channel study [40] by using the Weinberg-Tomozawa term within the traditional CUA, where the light-quark mass dependence of the decay constant and the adjustable subtraction constants are neglected. In contrast, we incorporates the pion mass dependence of all the model variables based on the CLS configuration. Avoid introducing the finite cutoff or the subtractive constants, we obtain the renormalized scattering amplitudes. In this sense that, our LO results of the quark mass dependence of $\Lambda(1405)$, illustrated in Fig. 4, is a model-independent prediction. The tendency of pole evolution could be checked by the forthcoming studies from the BaSc Collaboration.
Furthermore, it is worth mentioning that in Refs. [36, 39, 41], the evolution of two $\Lambda(1405)$ poles has been studied in the SU(3) trajectory. That’s corresponding to a trajectory from the SU(3) flavor limit to the physical point by keeping the average quark mass $\bar{m}=\frac{1}{3}(2m_{u/d}+m_{s})$ as a constant and varying a scaling parameter $x\in[0,1]$, which is introduced in the pseudoscalar meson masses, octet baryon masses, the meson decay constants, and the subtraction constants. At LO studies, the higher pole changes from a bound state to a resonance varying $x$ from 00 to $1$; correspondingly, the lower pole undergoes the bound state to a virtual state and finally comes to a resonance. However, in the recent NLO study [41], the evolutions of both $\Lambda(1405)$ poles along the SU(3) trajectory are changed a lot by including the NLO corrections. It gives us a hint that one needs to go a step further and include the NLO corrections in our framework, at least, to see the changes of the evolution results shown in Fig. 4. Such a study is beyond the scope of the present manuscript and will be reported in future publications.